# Router-Tuning: A Simple and Effective Approach for Dynamic Depth

### v4

@@ Abstract
The Mixture of Depths (MoD) was introduced to dynamically adjust the computational depth by skipping less important layers, thereby reducing computation and improving efficiency. Despite its promise, existing MoD approaches remain under-explored and face two main challenges: (1) high training costs due to the need to train the entire model along with the routers that determine which layers to skip, and (2) the risk of performance degradation when important layers are bypassed. In response to the first issue, we propose Router-Tuning, which fine-tunes only the routers on a small dataset, drastically reducing the computational overhead associated with full model training. For the second challenge, we investigate Router-Tuning across different architectures and granularities, demonstrating its effectiveness on Attention layers and MoE layers. This method preserves the model’s performance while significantly enhancing computational and memory efficiency. Extensive experiments demonstrate that our approach delivers competitive results while dramatically improving the computation efficiency, e.g., 21% speedup and only a 0.2% performance drop. The code is released at https://github.com/CASE-Lab-UMD/Router-Tuning.
@@ 1 Introduction
These challenges prompt us to consider two key questions: (1) How can we implement dynamic depth while minimizing training costs and time? (2) How can we enhance model efficiency without compromising performance, achieving faster inference while maintaining competitive task accuracy?
To tackle the first challenge, we introduce Router-Tuning, a novel method that fine-tunes only the router network without updating the backbone model’s parameters. As each router network is a lightweight, single-layer projector that accounts for less than 0.01% of the total parameters, the training overhead is minimal and even significantly lower than that of parameter-efficient finetuning methods Houlsby et al. (2019a); He et al. (2021) like LoRA Hu et al. (2022). Router-tuning requires only a small-scale dataset and fewer training steps, eliminating the need for large-scale pretraining or extensive continual training. This makes it a highly efficient and scalable solution.
To address the second challenge, we conduct a comprehensive investigation of target modules (e.g., Block, MLP, Attention, MoE), and various granularities (e.g., token and sequence). For dense transformer architectures, we propose Attention with Dynamic Depths, which selectively applies dynamic depth to attention layers. By focusing on attention layers known to exhibit high redundancy He et al. (2024b), Router-Tuning not only preserves model accuracy but also alleviates computational and memory bottlenecks. For MoE layers, which significantly impact efficiency due to the computational demands of multiple expert networks, we deploy Router-Tuning on each expert network to enhance the overall efficiency.
@@
  • We introduce Router-Tuning, a method that fine-tunes only the router on a small dataset, addressing the high computational cost of training the entire model with routers. This approach drastically reduces training overhead while maintaining effectiveness.
  • To explore the trade-off between performance and efficiency, we investigate target modules, deployment granularities, and diverse architectures, demonstrating the effectiveness of Router-Tuning on Attention and MoE layers.
  • Through comprehensive experiments, our approach demonstrates competitive results and significant efficiency improvements. Additionally, Router-Tuning can be seamlessly integrated with LoRA fine-tuning to further enhance both performance and efficiency.
@@ 2 Related Work
While increasing the depth of large language models has demonstrated promising performance across a wide range of tasks OpenAI et al. (2024); Team (2024), it also introduces layer redundancy Gromov et al. (2024); He et al. (2024b), posing efficiency challenges. To address this issue, several approaches have been proposed to reduce model depth Men et al. (2024) while maintaining comparable performance. Surprisingly, removing redundant layers has been shown to preserve performance while significantly reducing memory and computational costs Gromov et al. (2024); He et al. (2024a, b). Specifically, Gromov et al.Gromov et al. (2024) suggest dropping continuous Transformer blocks, and He et al. He et al. (2024b) propose fine-grained layer dropping to further improve the effectiveness of layer reduction. However, these static techniques fail to account for the varying complexity of different input sequences, where excessive layer removal can significantly degrade performance on more complex tasks. Instead of statically removing unimportant layers, our approach focuses on dynamically skipping these layers based on the specific inputs.
@@
Dynamic Depth, which allocates different layers based on the specific input, is an effective technique for accelerating inference while preserving performance Han et al. (2021, 2022). Recent works primarily implement dynamic depth through two key methods: Early-Exit Bae et al. (2023); Elhoushi et al. (2024) and Mixture of Depths (MoD) Raposo et al. (2024). Early-exit strategies terminate computation in later layers once sufficient confidence is achieved, effectively reducing redundant computations. In contrast, MoD offers greater flexibility by dynamically skipping less critical layers, enhancing adaptability and representational capacity. Despite their advantages, both Early-Exit and MoD often involve significant training overhead. For instance, LayerSkip Elhoushi et al. (2024) and MoD Raposo et al. (2024) require training models from scratch or extensive continual training, while Tan et al. Tan et al. (2024) extends pre-trained model training over long schedules to achieve optimal performance. To overcome these limitations, we propose Router-Tuning, an efficient approach to dynamic layer skipping that requires minimal additional offline training, providing a more practical and cost-effective solution.
@@ 3.1 Motivation
The Mixture of Depths (MoD) framework Raposo et al. (2024), which dynamically adjusts layer depth based on input complexity to enhance computational efficiency, was originally designed for integration during the pretraining phase, where transformer models are trained from scratch with MoD-enabled layers. More recently, Tan et al. (2024) applied MoD to pretrained Llama models Touvron et al. (2023) through continual training. While these approaches have demonstrated promising results, deploying MoD remains computationally expensive and time-consuming, posing challenges for scalability and real-world deployment. A more efficient alternative is to apply MoD directly to existing pretrained models, followed by lightweight fine-tuning of a subset of parameters Houlsby et al. (2019b); Hu et al. (2022), significantly reducing both computational costs and training time.
@@
These insights motivate us to deploy MoD in a cost-effective manner (i.e., Router-Tuning) while optimizing the balance between performance and efficiency.
@@ 3.2 Router-Tuning for Dynamic Depth
where $\bm{F}$ denotes a given layer and $\bm{y}$ is the output. This formulation ensures that the router is fully trainable through the following gradient calculations:
@@ 3.3 Extension to Mixture of Experts
where $\mathcal{K}$ denotes the indices of the top-$k$ selected experts, and $\bm{G}(\bm{x})_{i}$ represents the selection score for the $i$-th expert. By dynamically skipping experts within each layer, Router-Tuning significantly reduces both communication and computation costs.
@@ 3.4 Training Objectives
Given the computationally intensive nature of training entire LLMs and the constraints of real-world computational resources, our goal is to implement dynamic depth while minimizing both computational costs and time overhead. To achieve this, we focus exclusively on fine-tuning the router, as illustrated in Figure 1, thereby eliminating the need for costly full-model training.
@@ 4 Experiment Setup
We set $\tau$ as 0.5, which corresponds to the midpoint of the sigmoid function. To ensure that training starts from dense models, we initialize $\bm{W}$ to zero, ensuring that $\bm{R}(\bm{x})\geq\tau$ initially, i.e., training from dense models. To achieve the desired MoD capacity, we perform a grid search over the learning rate from {1e-5, 2e-5, 5e-5, 1e-4, 2e-4} and the scale factor $\lambda$ from {0, 0.1, 0.01, 0.001}, respectively
Table 2: Experimental results of Router-Tuning deployed at different granularities, assessing both task performance and inference efficiency. While Router-Tuning is primarily applied to Attention layers, we also evaluate its performance on Block and MLP layers for comparison. The number of skippable layers is constrained to 16, and the overall MoD capacity is 50%.
@@
$\mathrm{MLP}_{token}$ 1.05×1.05$\times$1.05 × 45.3 77.8 65.1 62.8 33.7 71.9 66.8 72.4 62.0
$\mathrm{MLP}_{seq}$ 1.06×1.06$\times$1.06 × 45.1 77.7 65.4 62.4 33.4 71.6 66.4 72.1 61.8
@@
$\mathrm{MLP}_{token}$ 1.05×1.05$\times$1.05 × 41.4 74.9 59.3 64.8 31.6 67.8 66.4 68.4 59.3
$\mathrm{MLP}_{seq}$ 1.06×1.06$\times$1.06 × 41.8 75.1 59.3 64.5 31.2 68.2 66.7 68.8 59.5
@@ 5.1 Performance of Router-Tuning
Router-Tuning achieves superior performance in Attention layers We first compare deploying Router-Tuning to different modules, e.g., Block, MLP, and Attention, as shown in Table 2. Based on the observation that deeper layers are more redundant than shallow layers Gromov et al. (2024); He et al. (2024b), we focus on deploying Router-Tuning to the deepest layers except the last one, leaving other layers unchanged.
@@
Figure 2: Comparison with Attention Drop under identical skipping ratios, highlighting the performance benefits of Router-Tuning.
Figure 3: Layer-wise Skipping Ratios for MoD and Attention Drop under overall skipping ratios of $50\%$.
Router-Tuning significantly outperforms static layer dropping
While Attention Drop He et al. (2024b), which statically drops attention layers, has demonstrated the promising performance of Layer Drop, its static nature lacks flexibility and limits representational power. Here, we further investigate the improvements offered by the dynamic mechanism. As shown in Figure 2, we compare Router-Tuning with Attention Drop He et al. (2024b) under the same computation budget — for instance, dropping 4 layers versus deploying MoD to 8 layers with 50% capacity. Under the same skipping ratios, Router-Tuning significantly outperforms Attention Drop He et al. (2024b) on the GSM8K benchmark Cobbe et al. (2021), e.g. 6.5% when the skipping ratio is 25.0%. In Figure 3, we further visualize the layer-wise skipping ratios of MoD versus Attention Drop. Unlike static approaches that permanently remove certain layers, Router-Tuning maintains the utilization of all layers by adaptively distributing skipping ratios across them. This flexible allocation strategy contributes to improved performance.
Table 3: Ablation study of Router-Tuning across different models architectures, including Llama Touvron et al. (2023), Qwen Bai et al. (2023), Mistral Jiang et al. (2023).
Llama-2-13B 1.00×1.$0$ 45.2 80.5 65.0 76.2 80.7 59.4 82.2 54.6 68.0
Qwen-2.5-14B 1.00×1.$0$ 45.6 82.2 79.1 80.4 85.3 67.2 84.3 79.7 75.5
Qwen-2.5-7B 1.00×1.$0$ 47.2 79.6 81.2 76.6 84.6 63.7 80.2 74.1 73.4
Mistral-7B 1.00×1.$0$ 44.4 82.2 68.2 79.0 82.2 60.6 83.2 62.4 70.3
@@ 5.2 Efficiency Improvements
The training efficiency of our method lies in two perspectives: trainable parameters and training steps. Since the router projects the input from dimension $d$ to 1, the number of trainable parameters is $d\times 1$ per layer, and the total number of trainable parameters is fewer than 0.01% of the whole model. Additionally, Router-Tuning only requires a few steps, which is verified in Section 6. Consequently, Router-Tuning on a single Nvidia A6000 GPU only takes less than 15 minutes, which is over 1000 times faster than DLO Tan et al. (2024) that finetunes the entire model and requires 36 hours on Nvidia RTX A100 GPUs.
@@
We also evaluate the run-time speed improvements achieved with Router-Tuning. The inference speed is measured throughout the entire generation process, including prefilling and generation. To ensure that our results accurately reflect the performance gains, we adhere to two key principles: (1) all operations are performed on a single Nvidia RTX A6000 Ada GPU, eliminating any communication overhead from multi-GPU setups; and (2) we increase batch sizes to fully utilize the GPU for each model. As shown in Table 2, different granularities contribute to different levels of speedup. This variation is primarily due to attention layers, where fine-grained token-level MoD introduces differing token lengths within a batch, necessitating padding operations to standardize sequence lengths. Instead, the speedups at the sequence level surpass those at the token level, with Router-Tuning achieving a 21% improvement in inference speed.
@@
Figure 4: Ablation study on the number of MoD layers.

### v6

@@ Abstract
The Mixture of Depths (MoD) was introduced to improve computational efficiency by dynamically skipping less important layers, reducing redundant computation while maintaining model capacity. Despite its promise, existing MoD approaches remain under-explored and face two main challenges: (1) high training costs due to the need to train the entire model along with the routers that determine which layers to skip, and (2) performance degradation when important layers are bypassed. In response to the first issue, we propose Router-Tuning, which fine-tunes only the routers on a small dataset, drastically reducing the computational overhead associated with full model training. For the second challenge, we investigate Router-Tuning across different architectures and granularities, demonstrating its effectiveness on Attention layers and MoE layers. This method preserves the model’s performance while significantly enhancing computational and memory efficiency. Extensive experiments demonstrate that our approach delivers competitive results while dramatically improving the computation efficiency, e.g., 21% speedup and only a 0.2% performance drop. The code is released at https://github.com/CASE-Lab-UMD/Router-Tuning.
@@ 1 Introduction
These challenges prompt us to reflect on the two key questions: (1) How can we implement dynamic depth to improve efficiency without incurring excessive training costs? (2) How can we preserve model performance in the presence of dynamic depth?
To tackle the first challenge, we introduce Router-Tuning, a novel method that fine-tunes only the router network without updating the backbone model’s parameters. As each router network is a lightweight, single-layer projector that accounts for less than 0.01% of the total parameters, the training overhead is minimal and even significantly lower than that of parameter-efficient finetuning methods Houlsby et al. (2019a); He et al. (2021) like LoRA Hu et al. (2022). Router-tuning requires only a small-scale dataset and fewer training steps, eliminating the need for large-scale pretraining or extensive continual training. Meanwhile, by freezing the backbone, Router-Tuning contributes to mitigating catastrophic forgetting and better retaining the original model performance Houlsby et al. (2019b); Liu et al. (2024); Qiao and Mahdavi (2024). These properties make Router-Tuning a highly efficient and scalable solution for dynamic adaptation.
To address the second challenge, we conduct a comprehensive investigation of target modules (e.g., Block, MLP, Attention, MoE), and various granularities (e.g., token and sequence). For dense transformer architectures, we propose Attention with Dynamic Depths, which selectively applies dynamic depth to attention layers. By focusing on attention layers known to exhibit high redundancy He et al. (2024b), Router-Tuning not only preserves model accuracy but also alleviates computational and memory bottlenecks. In the case of Mixture-of-Experts (MoE) layers Shazeer et al. (2017); Fedus et al. (2022), where efficiency is often hindered by the computational cost of activating multiple expert networks, we apply Router-Tuning at the expert level to enhance overall efficiency.
@@
  • We introduce Router-Tuning, a lightweight method that fine-tunes only the router using a small dataset, effectively addressing the high computational cost of training the entire model with routers.
  • We systematically investigate routing scopes, deployment granularities, and model architectures, demonstrating the effectiveness of Router-Tuning on Attention and MoE layers.
  • Through comprehensive experiments, Router-Tuning achieves competitive performance while delivering substantial improvements in training and inference efficiency.
@@ 2 Related Work
While increasing the depth of large language models has demonstrated promising performance across a wide range of tasks OpenAI et al. (2024); Team (2024), it also introduces layer redundancy Gromov et al. (2024); He et al. (2024b), posing efficiency challenges. To address this issue, several approaches have been proposed to reduce model depth Men et al. (2024) while maintaining comparable performance. Surprisingly, removing redundant layers has been shown to preserve performance while significantly reducing memory and computational costs Gromov et al. (2024); He et al. (2024a, b). Specifically, Gromov et al. (2024) suggest dropping continuous Transformer blocks, and He et al. (2024b) propose fine-grained layer dropping to further improve the effectiveness of layer reduction. However, these static techniques fail to account for the varying complexity of different input sequences, where excessive layer removal can significantly degrade performance on more complex tasks. Instead of statically removing unimportant layers, our approach focuses on dynamically skipping these layers based on the specific inputs.
@@
Dynamic Depth, which allocates different layers based on the specific input, is an effective technique for accelerating inference while preserving performance Han et al. (2021, 2022). Recent works primarily implement dynamic depth through two key methods: Early-Exit Bae et al. (2023); Elhoushi et al. (2024) and Mixture of Depths (MoD) Raposo et al. (2024). Early-exit strategies terminate computation in later layers once sufficient confidence is achieved, effectively reducing redundant computations. In contrast, MoD offers greater flexibility by dynamically skipping less critical layers, enhancing adaptability and representational capacity. Despite their advantages, both Early-Exit and MoD often involve significant training overhead. For instance, LayerSkip Elhoushi et al. (2024) and MoD Raposo et al. (2024) require training models from scratch or extensive continual training, while Tan et al. Tan et al. (2024) extends pre-trained model training over long schedules to achieve optimal performance. To overcome these limitations, we propose Router-Tuning, an efficient approach to dynamic layer skipping that requires minimal additional offline training, providing a more cost-effective solution.
@@ 3.1 Motivation
The Mixture of Depths (MoD) framework Raposo et al. (2024), which dynamically adjusts layer depth based on input complexity to enhance computational efficiency, was originally designed for integration during the pretraining phase, where transformer models are trained from scratch with MoD-enabled layers. More recently, Tan et al. (2024) applied MoD to pretrained Llama models Touvron et al. (2023) through continual training. While these approaches have demonstrated promising results, training with MoD remains computationally expensive and time-consuming, posing challenges for scalability and real-world deployment. A more efficient alternative is to apply MoD directly to existing pretrained models, followed by lightweight fine-tuning of a subset of parameters Houlsby et al. (2019b); Hu et al. (2022), significantly reducing both computational costs and training time.
@@
Building on these insights, we propose Router-Tuning, a cost-effective formulation of MoD that achieves a favorable trade-off between performance and computational costs.
@@ 3.2 Router-Tuning for Dynamic Depth
where $\bm{F}$ denotes a given layer and $\bm{y}$ is the output. This formulation ensures that the router is fully trainable through the gradient calculations:
@@ 3.3 Extension to Mixture of Experts
where $\mathcal{K}$ denotes the indices of the top-$k$ selected experts, and $\bm{G}(\bm{x})_{i}$ represents the selection score for the $i$-th expert. By dynamically skipping experts within each layer, Router-Tuning significantly reduces computation costs.
@@ 3.4 Training Objectives
Given the computationally intensive nature of training entire LLMs and the constraints of real-world computational resources, our goal is to implement dynamic depth while minimizing both computational costs and time overhead. To achieve this, we focus exclusively on fine-tuning the routers, as illustrated in Figure 1, thereby eliminating the need for costly full-model training.
@@ 4 Experiment Setup
We set $\tau$ as 0.5, which corresponds to the midpoint of the sigmoid function. To ensure that training starts from dense models, we initialize $\bm{W}$ to zero, ensuring that $\bm{R}(\bm{x})\geq\tau$ initially, i.e., training from dense models. To achieve the desired MoD capacity, we perform a grid search over the learning rate from {1e-5, 2e-5, 5e-5, 1e-4, 2e-4} and the scale factor $\lambda$ from {0, 0.1, 0.01, 0.001}, respectively.
Table 2: Router-Tuning at different granularities. We compare deployments on Attention, Block, and MLP layers. The number of skippable layers is capped at 16, with 50% MoD capacity. SpeedUp denotes inference-time speedup.
@@
$\mathrm{MLP}_{token}$ $1.05\times$ 45.3 77.8 65.1 62.8 33.7 71.9 66.8 72.4 62.0
$\mathrm{MLP}_{seq}$ $1.06\times$ 45.1 77.7 65.4 62.4 33.4 71.6 66.4 72.1 61.8
@@
$\mathrm{MLP}_{token}$ $1.05\times$ 41.4 74.9 59.3 64.8 31.6 67.8 66.4 68.4 59.3
$\mathrm{MLP}_{seq}$ $1.06\times$ 41.8 75.1 59.3 64.5 31.2 68.2 66.7 68.8 59.5
@@ 5.1 Performance of Router-Tuning
Router-Tuning achieves superior performance on Attention layers We first compare deploying Router-Tuning to different modules, e.g., Block, MLP, and Attention, as shown in Table 2. Based on the observation that deeper layers are more redundant than shallow layers Gromov et al. (2024); He et al. (2024b), we focus on deploying Router-Tuning to the deepest layers except the last one, leaving other layers unchanged.
@@
Figure 2: Comparison between Router-Tuning and Layer Drop on MLP and Attention layers under a fixed 25% overall skipping ratio.
Router-Tuning improves over static layer dropping
While statically dropping attention layers He et al. (2024b) has demonstrated promising performance, its static nature lacks flexibility and limits representational power. Here, we further investigate the improvements offered by the dynamic mechanism. Figure 2 compares Router-Tuning with static Layer Drop He et al. (2024b), where Router-Tuning consistently achieves superior performance. For more complex tasks that are more sensitive to layer skipping, as shown in Figure 3, we compare Router-Tuning with Layer Drop on attention layers (i.e., “Attention Drop”) under the same computation budget, e.g., dropping 4 layers versus deploying MoD to 8 layers with 50% capacity. Under the same skipping ratios, Router-Tuning significantly outperforms Attention Drop on the GSM8K benchmark Cobbe et al. (2021), e.g., 6.5% when the skipping ratio is 25.0%. In Figure 4, we further visualize the layer-wise skipping ratios of MoD versus Attention Drop. Unlike static approaches that permanently remove certain layers, Router-Tuning maintains the utilization of all layers by adaptively distributing skipping ratios across them. This flexible allocation strategy contributes to improved performance.
Figure 3: Comparison with Attention Drop on GSM8K tasks under identical skipping ratios.
Figure 4: Layer-wise Skipping Ratios for Attention layers after Layer Drop and Router-Tuning.
Router-Tuning outperforms dynamic skipping methods
Table 3 compares Router-Tuning with dynamic skipping baselines, including DLO Tan et al. (2024) and Skip Transformer Peroni and Bertsimas (2024). Following the setup of baselines, we conduct the comparison on the token level for MLP layers. Router-Tuning consistently outperforms these methods across most tasks, despite operating under the router-only training constraint. On the one hand, Router-Tuning freezes the model backbones, which contributes to avoiding the risks of catastrophic forgetting Houlsby et al. (2019b); Liu et al. (2024); Qiao and Mahdavi (2024). Additionally, Router-Tuning is trained end-to-end without being constrained by precomputed labels (e.g., token-level similarity scores in DLO) or stochastic gating mechanisms. This design enables more flexible and stable learning, while preserving the pretrained capabilities of the backbone.
Table 3: Performance comparison against dynamic dropping baselines, including DLO Tan et al. (2024) and Skip Transformer Peroni and Bertsimas (2024).
Method ARC-C HellaSwag MMLU WinoG Avg.
DLO 44.5 64.2 62.1 71.3 60.5
Skip Transformer 44.7 64.4 62.4 71.5 60.8
Router-Tuning 45.3 65.1 62.8 72.4 61.4
@@ 5.2 Efficiency Improvements
Table 4: Comparison of training strategies of achieving dynamic layer skipping. The training time for MoD is left blank as it was not conducted on LLaMA-3-8B.
Method Target Modules Granularity Training Stage Trainable Training Time
MoD Raposo et al. (2024) Block Token Pretraining Full Model
DLO Tan et al. (2024) MLP Token Continual Pretraining Full Model 36h on NVIDIA A100
Router-Tuning Block / MLP / Attn Token / Sequence Finetuning Router 15m on NVIDIA A6000
@@
The training efficiency of our method lies in two perspectives: trainable parameters and training steps. Since the router projects the input from dimension $d$ to 1, the number of trainable parameters is $d\times 1$ per layer, and the total number of trainable parameters is fewer than 0.01% of the whole model. Additionally, Router-Tuning only requires a few steps, which is verified in Section 6. Consequently, as shown in Table 4, Router-Tuning can be completed in under 15 minutes on a single NVIDIA A6000 GPU—over 1000 times faster than DLO Tan et al. (2024), which performs large-scale training on full models and takes 36 hours on NVIDIA RTX A100 GPUs.
@@
We also evaluate the run-time speed improvements achieved with Router-Tuning. The inference speed is measured throughout the entire generation process, including prefilling and generation. To ensure that our results accurately reflect the performance gains, we adhere to two key principles: (1) all operations are performed on a single Nvidia RTX A6000 Ada GPU, eliminating any communication overhead from multi-GPU setups; and (2) we set the maximal sequence length as 2048 and increase batch sizes to fully utilize the GPU for each model.
As shown in Table 2, skipping attention layers yields a more substantial speedup than skipping MLP layers, which is primarily due to the quadratic complexity of the attention mechanism and the memory overhead associated with KV-cache Zhang et al. (2023); Singhania et al. (2024); He et al. (2024b). On the other hand, different granularities contribute to different levels of speedup. This variation is primarily due to attention layers, where fine-grained token-level MoD introduces differing token lengths within a batch, necessitating padding operations to standardize sequence lengths. Instead, the speedups at the sequence level surpass those at the token level, with Router-Tuning achieving a 21% improvement in inference speed. Therefore, we set Router-Tuning on the sequence level for attention layers as the default setting.
@@
Table 5: Ablation study of Router-Tuning across multiple model architectures and scales, highlighting its robustness and consistent improvements.
Llama-2-13B $1.00\times$ 45.2 80.5 65.0 76.2 80.7 59.4 82.2 54.6 68.0
Qwen-2.5-14B $1.00\times$ 45.6 82.2 79.1 80.4 85.3 67.2 84.3 79.7 75.5
Qwen-2.5-7B $1.00\times$ 47.2 79.6 81.2 76.6 84.6 63.7 80.2 74.1 73.4
Mistral-7B $1.00\times$ 44.4 82.2 68.2 79.0 82.2 60.6 83.2 62.4 70.3